# NeuralFuse: Learning to Recover the Accuracy of Access-Limited Neural Network Inference in Low-Voltage Regimes

### v2

@@ 4 Experiments
In this section, we present the experimental setup and results of NeuralFuse on different datasets and various architectures. In addition, we also provide the visualization results, detailed analysis, and ablation studies to better understand the properties of NeuralFuse.
@@ 4.1 Experiment Setups
We evaluate NeuralFuse on four different datasets: CIFAR-10 (Krizhevsky & Hinton, 2009), CIFAR-100 (Krizhevsky & Hinton, 2009), GTSRB (Stallkamp et al., 2012), and ImageNet-10 (Deng et al., 2009). CIFAR-10 consists of ten classes, with 50,000 training images and 10,000 for testing. Similarly, CIFAR-100 consists of 100 classes, with 500 training images and 100 testing images in each class. GTSRB (German Traffic Sign Recognition Benchmark) is a dataset that contains 43 classes with 39,209 training images and 12,630 testing images. Similar to CIFAR-10 and CIFAR-100, we resize GTSRB into 32323 in our experiment. For ImageNet-10, we chose the same ten categories as Huang et al. (2022), and there are 13,000 training images and 500 test images, which are cropped into 2242243. Due to the space limit, the CIFAR-100 results are given in Appendix F and G.
Base Models.
@@
We select several common architectures for our base models: ResNet18, ResNet50 (He et al., 2016), VGG11, VGG16, and VGG19 (Simonyan & Zisserman, 2015). In order to meet the setting of deploying models on a chip, all of our based models are trained by quantization-aware training following Stutz et al. (2021).
NeuralFuse Generators.
@@
The architecture of the NeuralFuse generator () is based on the Encoder-Decoder structure. We design and compare three types of generators, namely Convolution-based, Deconvolution-based, and UNet-based. We also consider large(L)/small(S) network sizes for each type. Both Convolution-based and Deconvolution-based variants follow a similar architecture for ease of comparison. Also, the generators were trained with quantization-aware training. More details are given in Appendix B.
@@
Convolution-based (Conv). Conv uses Convolution with MaxPool layers for the encoder and Convolution with UpSample layers for the decoder. This architecture has shown its efficiency and effectiveness for generating input-aware backdoor triggers (Nguyen & Tran, 2020).
@@
Deconvolution-based (DeConv). DeConv uses Convolution with MaxPool layers for the encoder and Deconvolution layers for the decoder. This modification is expected to enhance its performance and reduce energy consumption.
@@
UNet-based (UNet). UNet uses Convolution with MaxPool layers for the encoder and Deconvolution layers for the decoder. UNet is known for its robust performance in image segmentation (Ronneberger et al., 2015), and thus, we incorporated it as one of our implementations.
Energy Consumption Calculation.
@@
In this paper, we focus on the resilience to low-voltage-induced bit errors in model weights, and our reported energy consumption in Figure 1 is based on the product of the total number of SRAM memory accesses in a systolic-array-based CNN accelerator and the dynamic energy per read access at a given voltage. Research by Chen et al. (2016) also shows that the energy consumption of the SRAM part (buffer and array) accounts for a large part of the total system energy consumption. We assume that there is no bit error on NeuralFuse, as we consider it as an add-on data preprocessing module that could also be performed by the general-purpose core. Therefore, when implemented in the accelerator, NeuralFuse computation is performed at a higher error-free voltage with a dynamically scaled voltage supply to the memories. We report the reduction in overall weight memory energy consumption (i.e., NeuralFuse + Base Model under bit error rate) with respect to the unprotected base model in the regular voltage mode (i.e., bit error rate and without NeuralFuse).
@@
To obtain the number of memory accesses, we used the SCALE-SIM simulator (Samajdar et al., 2020), and our chosen configuration simulates an output-stationary dataflow and a 3232 systolic array with 256KB of weight memory. We obtained the dynamic energy per read access of the SRAM at the minimum voltage () and at the voltage corresponding to a bit error rate () from Cadence ADE Spectre simulations, both at the same clock frequency.
Relaxed and Restricted Access Settings.
@@
Our experiments consider two scenarios (relaxed/restricted access). For relaxed access, the information of the base model is not entirely transparent but it allows obtaining gradients from the black-box model through backpropagation. Therefore, this setting allows direct training of NeuralFuse with the base model using EOPM. On the other hand, for restricted access, only the inference function is allowed for the base model. Therefore, we train NeuralFuse by using a white-box surrogate base model and then transfer the generator to the access-restricted model.
Computing Resources.
@@
Our experiments are performed using 8 Nvidia Tesla V100 GPUs and are implemented with PyTorch. NeuralFuse generally takes 150 epochs to converge, and its training time is similar for the same base model. On both CIFAR-10 and CIFAR-100 datasets, the average training times were 17 hours (ResNet18), 50 hours (ResNet50), 9 hours (VGG11), 13 hours (VGG16), and 15 hours (VGG19). For GTSRB, the average training times were 9 hours (ResNet18), 27 hours (ResNet50), 5 hours (VGG11), 7 hours (VGG16), and 8 hours (VGG19). For ImageNet-10, the average training times were 32 hours (ResNet18), 54 hours (ResNet50), 50 hours (VGG11), 90 hours (VGG16), and 102 hours (VGG19).
(a) CIFAR-10, Bit Error Rate
(b) GTSRB, Bit Error Rate
(c) ImageNet-10, Bit Error Rate
4.2 Performance Evaluation on Relaxed Access Setting
@@
The experimental results of the Relaxed Access setting are shown in Figure 3. We train and test NeuralFuse with various base models (ResNet18, ResNet50, VGG11, VGG16, and VGG19). Two power settings have been considered: nominal voltage (no bit error) and low voltage (random bit errors), and the corresponding bit error rate (B.E.R.) due to low voltage is (CIFAR-10, GTSRB) and (ImageNet-10). The B.E.R. of ImageNet-10 is lower because the pre-trained models have more parameters than CIFAR-10 and GTSRB. For each experiment, we sample perturbed models (independent from training) for evaluation and report the mean and standard deviation of the test accuracy. In the following, clean accuracy (C.A.) means that the model is measured at nominal voltage, and perturbed accuracy (P.A.) means that it is measured at low voltage.
@@
For CIFAR-10 and GTSRB, we observe that large generators like ConvL and UNetL can significantly recover the perturbed accuracy in the range of to on ResNet18, VGG11, VGG16, and VGG19. For ResNet50, the recover percentage is slightly worse than other base models, but it can attain up to recover percentage on GTSRB. On the other hand, the recover percentage based on small generators like DeConvS are worse than larger generators. This can be explained by the better ability to learn error-resistant generators for larger-sized networks (though they may consume more energy). For ImageNet-10, using larger generators can also gain better performance recovery on perturbed accuracy. This also demonstrates that NeuralFuse can still work well even with large input sizes and is applicable to different datasets.
4.3 Transferability for Restricted Access Setting
@@
In the Restricted Access scenario, we train NeuralFuse generators on a white-box surrogate base model and transfer it to other black-box base models. The experimental results are shown in Table 1. We adopt ResNet18 and VGG19 as the white-box surrogate (source) models for training the generators under a bit error rate. For the generators, we choose ConvL and UNetL as they obtain better performance in Figure 3.
From Table 1, we can find that transferring from a larger B.E.R. like can give strong resilience to a smaller B.E.R. like or . We also find that using VGG19 as a surrogate model with UNet-based generators like UNetL can give better recovery than other combinations. On the other hand, in some cases, we observe that if we transfer between the same source and target models (but with different B.E.R. for training and testing), the performance may outperform the original relaxed-access results. For instance, when transferring VGG19 with UNetL under a B.E.R. to VGG19 or VGG11 under a B.E.R., the accuracy is v.s. for VGG19 (original), and v.s. for VGG11 (original), respectively. We conjecture that the generators trained on a larger B.E.R. can actually cover the error patterns of a smaller B.E.R., and even help improve generalization under a smaller B.E.R. These findings show great promise for recovering the accuracy of access-limited base models in low-voltage settings.
@@
S.M. T.M. B.E.R. C.A. P.A. ConvL (1.5%) UNetL (1.5%) C.A. (NF) P.A. (NF) R.P. C.A. (NF) P.A. (NF) R.P. ResNet18 ResNet18 1% 92.6 38.9 12.4 89.8 89.0 0.5 50.1 85.8 85.2 0.5 46.3 0.5% 70.1 11.6 89.6 0.2 19.5 85.7 0.2 15.6 ResNet50 1% 92.6 26.1 9.4 89.2 36.1 18 10.0 84.4 38.9 16 12.8 0.5% 61.0 10.3 74.1 10 13.1 72.7 4.6 11.7 VGG11 1% 88.4 42.2 11.6 86.3 59.2 10 17.0 82.3 69.8 7.5 27.6 0.5% 63.6 9.3 78.9 4.9 15.3 77.0 4.0 13.4 VGG16 1% 90.3 35.7 7.9 89.4 62.2 18 26.5 84.7 68.9 14 33.2 0.5% 66.6 8.1 83.4 5.5 16.8 80.5 5.9 13.9 VGG19 1% 90.5 36.0 12.0 89.8 49.9 23 13.9 85.0 55.1 17 19.1 0.5% 64.2 12.4 81.8 8.5 17.6 78.5 6.8 14.3 VGG19 ResNet18 1% 92.6 38.9 12.4 88.9 62.6 13 23.7 85.0 72.3 11 33.4 0.5% 70.1 11.6 84.2 7.2 14.1 82.1 2.2 12.0 ResNet50 1% 92.6 26.1 9.4 88.8 37.9 18 11.8 85.2 46.7 17 20.6 0.5% 61.0 10.3 76.6 7.8 15.6 78.3 3.7 17.3 VGG11 1% 88.4 42.2 11.6 88.9 76.0 6.1 33.8 85.5 81.9 3.9 39.7 0.5% 63.6 9.3 85.9 2.6 22.3 84.8 0.5 21.2 VGG16 1% 90.3 35.7 7.9 89.0 76.5 9.0 40.8 85.9 79.2 7.5 43.5 0.5% 66.6 8.1 87.7 0.7 21.1 84.7 0.9 18.1 VGG19 1% 90.5 36.0 12.0 89.1 80.2 12 44.2 86.3 84.3 1.2 48.3 0.5% 64.2 12.4 88.8 0.4 24.6 85.9 0.3 21.7 • [Note] S.M.: source model, used for training generators, T.M.: target model, used for testing generators, B.E.R.: the bit error rate of the target model, C.A. (): clean accuracy, P.A. (): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.
Base Model ConvL ConvS DeConvL DeConvS UNetL UNetS ResNet18 19.0 29.1 21.2 27.5 24.0 28.9 ResNet50 25.4 29.9 26.4 29.2 27.7 29.9 VGG11 6.6 27.5 11.2 24.1 17.1 27.2 VGG16 17.1 28.9 19.7 27.0 23.0 28.7 VGG19 20.3 29.7 22.3 27.8 24.8 29.1
4.4 Energy-Accuracy Tradeoff
@@
We report the total dynamic energy consumption as the total number of SRAM accesses times the dynamic energy of a single SRAM access. Specifically, we used SCALE-SIM to calculate the total weight memory access (T.W.M.A.), which can be found in Table 6 in Appendix C. In Table 2, we report the percentage of energy savings at voltages that would yield a bit error rate for various base model and generator combinations. The formula of the energy saving percentage (ES, ) is defined as:
where NV denotes nominal voltage, and LV denotes low voltage regime.
@@
The results demonstrate that when utilizing ResNet18 as a base model, NeuralFuse can recover model accuracy by while saving energy. In Appendix C, we provided more results on the tradeoff between energy and accuracy of different NeuralFuse and Base Model combinations. Overall, our results show promising results that using NeuralFuse can effectively restore model accuracy when the SRAM encounters low-voltage-induced random bit errors.
Runtime and Latency.
@@
On the other hand, runtime and its corresponding energy consumption may also affect the overall energy saving. For instance, previous research has shown that Multiply-and-accumulate (MAC) operations account for over 99% of total operations in state-of-the-art CNNs, dominating both processing runtime and energy consumption (Yang et al., 2017). Therefore, we also report the results of MACs-based energy consumption estimation in Appendix C, and latency report in Appendix D. We want to note that the additional latency overhead is an inevitable tradeoff for reducing energy consumption in our setting. Although runtime and latency are not the major focus of this paper, we envision future research designing a more lightweight version of the NeuralFuse module or using model compression techniques on NeuralFuse to reduce latency and runtime.
4.5 Model Size and Efficiency of NeuralFuse

### v3

@@ 4.1 Experiment Setups
Datasets. We evaluate NeuralFuse on four different datasets: CIFAR-10 (Krizhevsky and Hinton, 2009), CIFAR-100 (Krizhevsky and Hinton, 2009), the German Traffic Sign Recognition Benchmark (GTSRB) (Stallkamp et al., 2012), and ImageNet-10 (Deng et al., 2009). CIFAR-10 consists of 10 classes, with 50,000 training images and 10,000 testing images in total. Similarly, CIFAR-100 consists of 100 classes, with 500 training images and 100 testing images in each. The GTSRB contains 43 classes with a total of 39,209 training images and 12,630 testing images. Similar to CIFAR-10 and CIFAR-100, we resize GTSRB into 32323 in our experiment. For ImageNet-10, we chose the same ten categories as Huang et al. (2022), in which there are 13,000 training images and 500 test images cropped into 2242243. Due to space limitations, our CIFAR-100 results are presented in Appendices F and G.
@@
Base Models. We selected several common architectures for our base models: ResNet18, ResNet50 (He et al., 2016), VGG11, VGG16, and VGG19 (Simonyan and Zisserman, 2015). To replicate the deployment of models on chips, all our based models were given quantization-aware training that followed Stutz et al. (2021).
@@
NeuralFuse Generators. The architecture of the NeuralFuse generator () ) is based on an encoder-decoder structure. We designed and compared three types of generators, namely convolution-based, deconvolution-based, and UNet-based. We also considered large(L)/small(S) network sizes for each type. Further details can be found below and in Appendix B.
@@
Convolution-based (Conv). Conv uses convolution with MaxPool layers for its encoder and convolution with UpSample layers for its decoder. This architecture has previously been shown to be efficient and effective at generating input-aware backdoor triggers (Nguyen and Tran, 2020).
@@
Deconvolution-based (DeConv). DeConv uses convolution with MaxPool layers for its encoder and deconvolution layers for its decoder. We expected this modification both to enhance its performance and to reduce its energy consumption.
@@
UNet-based (UNet). UNet uses convolution with MaxPool layers for its encoder, and deconvolution layers for its decoder. UNet is known for its robust performance in image segmentation (Ronneberger et al., 2015).
@@
Energy-consumption Calculation. The energy consumption reported in Figure 1 is based on the product of the total number of SRAM memory accesses in a systolic array-based convolution neural network (CNN) accelerator and the dynamic energy per read access at a given voltage. Research by Chen et al. (2016) previously showed that energy consumption by SRAM buffers and arrays accounts for a high proportion of total system energy consumption. We assume that there are no bit errors on NeuralFuse, given its status as an add-on data preprocessing module whose functions could also be performed by a general-purpose core. In this work we assume it is implemented on the accelerator equipped with dynamic voltage scaling and therefore NeuralFuse computation is performed at nominal error-free voltage. We report a reduction in overall weight-memory energy consumption (i.e., NeuralFuse + Base Model under a bit-error rate) with respect to the unprotected base model in the regular-voltage mode (i.e., bit-error rate and without NeuralFuse).
@@
To quantify memory accesses, we used the SCALE-SIM simulator (Samajdar et al., 2020), and our chosen configuration simulated an output-stationary dataflow and a 3232 systolic array with 256KB of weight memory. We collected data on the dynamic energy per read access of the SRAM both at and at the voltage corresponding to a bit-error rate () from Cadence ADE Spectre simulations, both at the same clock frequency.
@@
Relaxed and Restricted Access Settings. In the first of our experiments’ two scenarios, relaxed access, the base-model information was not entirely transparent, but allowed us to obtain gradients from the black-box model through backpropagation. Therefore, this scenario allowed direct training of NeuralFuse with the base model using EOPM. In the restricted-access scenario, on the other hand, only the inference function was allowed for the base model, and we therefore trained NeuralFuse using a white-box surrogate base model and then transfering the generator to the access-restricted model.
(a) CIFAR-10, Bit-error Rate
(b) GTSRB, Bit-error Rate
(c) ImageNet-10, Bit-error Rate
@@
Computing Resources. Our experiments were performed using eight Nvidia Tesla V100 GPUs and implemented with PyTorch. NeuralFuse was found to generally take 150 epochs to converge, and its training time was similar to that of the base model it incorporated. On both the CIFAR-10 and CIFAR-100 datasets, average training times were 17 hours (ResNet18), 50 hours (ResNet50), 9 hours (VGG11), 13 hours (VGG16), and 15 hours (VGG19). For GTSRB, the average training times were 9 hours (ResNet18), 27 hours (ResNet50), 5 hours (VGG11), 7 hours (VGG16), and 8 hours (VGG19); and for ImageNet-10, the average training times were 32 hours (ResNet18), 54 hours (ResNet50), 50 hours (VGG11), 90 hours (VGG16), and 102 hours (VGG19).
4.2 Performance Evaluation, Relaxed-access Scenario
@@
Our experimental results pertaining to the relaxed-access scenario are shown in Figure 3. The bit-error rate (BER) due to low voltage was in the cases of CIFAR-10 and GTSRB, and for ImageNet-10. The BER of ImageNet-10 was lower than that of the other two because, being pre-trained, it has more parameters than either of them. For each experiment, we sampled and evaluated perturbed models (independent from training), and below, we report the means and standard deviations of their respective accuracies. Below, clean accuracy (CA) refers to a model’s accuracy measured at nominal voltage, and perturbed accuracy (PA) to its accuracy measured at low voltage.
@@
In the cases of CIFAR-10 and the GTSRB, we observed that large generators like ConvL and UNetL recovered PA considerably, i.e., in the range of to on ResNet18, VGG11, VGG16, and VGG19. ResNet50’s recovery percentage was slightly worse than those of the other base models, but it nevertheless attained up to recovery on the GTSRB. On the other hand, the recovery percentages achieved when we used small generators like DeConvS were worse than those of their larger counterparts. This could be explained by larger-sized networks’ better ability to learn error-resistant generators (though perhaps at the cost of higher energy consumption). In the case of ImageNet-10, using larger generators also yielded better PA performance recovery, further demonstrating NeuralFuse’s ability to work well with large input sizes and varied datasets.
4.3 Performance Evaluation, Restricted-access Scenario (Transferability)
@@
The experimental results of our restricted-access scenario are shown in Table 1. We adopted ResNet18 and VGG19 as our white-box surrogate source models for training the generators under a bit-error rate. We chose ConvL and UNetL as our generators because they performed best out of the six we tested (see Figure 3).
@@
SM TM BER CA PA ConvL (1.5%) UNetL (1.5%) CA (NF) PA (NF) RP CA (NF) PA (NF) RP ResNet18 ResNet18 1% 92.6 38.9 12.4 89.8 89.0 0.5 50.1 85.8 85.2 0.5 46.3 0.5% 70.1 11.6 89.6 0.2 19.5 85.7 0.2 15.6 ResNet50 1% 92.6 26.1 9.4 89.2 36.1 18 10.0 84.4 38.9 16 12.8 0.5% 61.0 10.3 74.1 10 13.1 72.7 4.6 11.7 VGG11 1% 88.4 42.2 11.6 86.3 59.2 10 17.0 82.3 69.8 7.5 27.6 0.5% 63.6 9.3 78.9 4.9 15.3 77.0 4.0 13.4 VGG16 1% 90.3 35.7 7.9 89.4 62.2 18 26.5 84.7 68.9 14 33.2 0.5% 66.6 8.1 83.4 5.5 16.8 80.5 5.9 13.9 VGG19 1% 90.5 36.0 12.0 89.8 49.9 23 13.9 85.0 55.1 17 19.1 0.5% 64.2 12.4 81.8 8.5 17.6 78.5 6.8 14.3 VGG19 ResNet18 1% 92.6 38.9 12.4 88.9 62.6 13 23.7 85.0 72.3 11 33.4 0.5% 70.1 11.6 84.2 7.2 14.1 82.1 2.2 12.0 ResNet50 1% 92.6 26.1 9.4 88.8 37.9 18 11.8 85.2 46.7 17 20.6 0.5% 61.0 10.3 76.6 7.8 15.6 78.3 3.7 17.3 VGG11 1% 88.4 42.2 11.6 88.9 76.0 6.1 33.8 85.5 81.9 3.9 39.7 0.5% 63.6 9.3 85.9 2.6 22.3 84.8 0.5 21.2 VGG16 1% 90.3 35.7 7.9 89.0 76.5 9.0 40.8 85.9 79.2 7.5 43.5 0.5% 66.6 8.1 87.7 0.7 21.1 84.7 0.9 18.1 VGG19 1% 90.5 36.0 12.0 89.1 80.2 12 44.2 86.3 84.3 1.2 48.3 0.5% 64.2 12.4 88.8 0.4 24.6 85.9 0.3 21.7 • Note. SM: source model, used for training generators; TM: target model, used for testing generators; BER: the bit-error rate of the target model; CA (): clean accuracy; PA (): perturbed accuracy; NF: NeuralFuse; and RP: total recovery percentage of PA (NF) vs. PA
From Table 1, we can see that transferring from a larger BER such as can endow a smaller one (e.g., ) with strong resilience. The table also makes it clear that using VGG19 as a surrogate model with UNet-based generators like UNetL can yield better recovery performance than other model/generator combinations. On the other hand, we observed in some cases that if we transferred between source and target models of the same type (but with different BERs for training and testing), performance results could exceed those we had obtained during the original relaxed-access scenario. For instance, when we transferred VGG19 with UNetL under a BER to VGG19 or VGG11 under a BER, the resulting accuracies were (as against for original VGG19) and (as against for original VGG11). We conjecture that generators trained on relatively large BERs can cover the error patterns of smaller BERs, and even help improve the latter’s generalization. These findings indicate the considerable promise of access-limited base models in low-voltage settings to recover accuracy.
ConvL ConvS DeConvL DeConvS UNetL UNetS ResNet18 19.0 29.1 21.2 27.5 24.0 28.9 ResNet50 25.4 29.9 26.4 29.2 27.7 29.9 VGG11 6.6 27.5 11.2 24.1 17.1 27.2 VGG16 17.1 28.9 19.7 27.0 23.0 28.7 VGG19 20.3 29.7 22.3 27.8 24.8 29.1
4.4 Energy/Accuracy Tradeoff
@@
We report total dynamic energy consumption as the total number of SRAM-access events multiplied by the dynamic energy of a single such event. Specifically, we used SCALE-SIM to calculate total weight-memory access (TWMA), specifics of which can be found in Appendix C’s Table 6. In Table 2, below, we report the percentages of energy saved (ES) at voltages that yield a bit-error rate for various base-model and generator combinations. The formula for computing ES is
where NV denotes nominal voltage regime, and LV, a low-voltage one.
@@
Our results indicate that when ResNet18 is utilized as a base model, NeuralFuse can recover model accuracy by while reducing energy use by . In Appendix C, we provide more results on the tradeoff between energy and accuracy of different NeuralFuse and base-model combinations. Overall, it would appear that using NeuralFuse can effectively restore model accuracy when SRAM encounters low-voltage-induced random bit errors.
@@
Runtime and Latency. On the other hand, runtime and its corresponding energy consumption may also affect overall energy savings. For instance, previous research has shown that multiply-and-accumulate (MAC) operations account for more than 99% of all operations in state-of-the-art CNNs, dominating processing runtime and energy consumption alike Yang et al. (2017). Therefore, we also report the results of our MAC-based energy-consumption estimation in Appendix C, and of our latency analysis in Appendix D. Here, it should also be noted that an additional latency overhead is an inevitable tradeoff for reducing energy consumption in our scenarios. Although neither runtime nor latency is a major focus of this paper, future researchers could usefully design a lighter-weight version of the NeuralFuse module, or apply model-compression techniques to it, to reduce these two factors.
Base Model BER NeuralFuse ConvL ConvS DeConvL DeConvS UNetL UNetS ResNet18 1% 67.5 182 76.6 190.7 94.5 245.9 0.5% 24.7 73.3 30.7 62.5 33.6 88.3 ResNet50 1% 37.4 75.1 57.4 102.7 102.3 248.4 0.5% 35.2 108.7 40.4 92.5 47.4 124.6 VGG11 1% 62.3 212.9 69.5 165.8 92.0 251.7 0.5% 32.3 96.3 35.8 77.2 38.9 100.7 VGG16 1% 69.6 211.2 76.9 196.5 98.8 292.9 0.5% 30.3 98.1 33.0 75.3 40.6 113 VGG19 1% 57.6 147.5 65.5 141.6 95.4 250.8 0.5% 33.0 91.0 37.5 70.2 43.1 106.4
4.5 Model Size and NeuralFuse Efficiency